OpenStudy (anonymous):

In the following sequence, the 1st yerm, the 2nd term and the 3rd term are 1,2 and 3 respectively. For any positive integer n, the (n+3)th term is the sum of the (n+2)th term, the (n+1)th term and the nth term. Find the 9th term of the sequence. 1, 2, 3, 6, 11, .....

OpenStudy (phi):

are you sure it goes 1,2,3 6 ?

OpenStudy (anonymous):

yes, the WorkSheet stated that

OpenStudy (phi):

ok, I mis-read it.

OpenStudy (phi):

add the last 3 numbers to get the next number

OpenStudy (cwrw238):

so next term = 11 + 6 + 3 = 20

OpenStudy (phi):

next one is 20= 3+ 6+ 11

OpenStudy (anonymous):

ohh yes

OpenStudy (anonymous):

and then 37?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

8th=68? 9th=125?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

oic